# SDP bounds on quantum codes

### v1

@@ 2.2. Proof sketch: Lovász and quantum Delsarte bounds
    We consider a relaxation of the complete hierarchy of Result B that is indexed by $\langle E_{a}\rangle E_{a}$ with $E_{a}\in\mathcal{E}_{n}$. The key object is a matrix with entries $\Gamma_{ab}=\Re({\langle E_{a}^{\dagger}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle})$. Such matrix is positive semidefinite by construction, and satisfies $\Gamma_{00}=1$, $\Gamma_{a0}=\Gamma_{aa}$, and $\Gamma_{ab}=0$ if $E_{a}E_{b}+E_{b}E_{a}=0$. For pure codes additionally $\Gamma_{ab}=0$ if $0<\operatorname{wt}(E_{a})<\delta$, $0<\operatorname{wt}(E_{b})<\delta$, or $0<\operatorname{wt}(E_{a}E^{\dagger}_{b})<\delta$.
@@
    Consider the case of self-dual codes ($K=1$). Then the set of matrices $\Gamma$ with this structure can be identified with the Lovász theta body for a graph with loops, that is the feasible region of Eq. (3.5),
@@
    Averaging $\mathrm{TH}(H)$ of the $4$-ary Hamming graph $H$ under distance preserving automorphisms $\operatorname{Aut}(4,n)$ yields, under an additional equality constraint, the quantum Delsarte bound [Eq. (3.4)].
@@
    Using the Terwilliger algebra for $4$-ary codes, we average this main relaxation over distance-preserving automorphisms that keep the zero codeword invariant $\operatorname{Aut}_{0}(4,n)$. A block-diagonalization using the Terwilliger algebra [23] yields a symmetry-reduced semidefinite programming bound of size $O(n^{4})$.
A variation of the Lovász bound recovers the non-existence of the $(\!(7,1,4)\!)_{2}$ code; while a variation of the symmetry-reduced semidefinite programming bounds shows the non-existence of $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$.
@@ 3. Key concepts
A quantum error-correcting code (QECC) with parameters $(\!(n,K,\delta)\!)_{2}$ encodes a Hilbert space of dimension $K$ in $n$ qubits. A QECC that can correct all errors acting on up to $\lfloor\tfrac{\delta-1}{2}\rfloor$ arbitrary qubit positions is said to have distance $\delta$.
Define a $n$-qubit error basis (or Pauli basis) ${\mathcal{E}_{n}}$ by considering all tensor products of the identity and the three Pauli matrices
(7) $\mathds{1}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\,,\quad X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,,\quad Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\,,\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\,.$
The elements of $\mathcal{E}_{n}$ then generate the $n$-qubit Pauli group $P_{n}=\{i^{b}E_{a}\,|\,E_{a}\in\mathcal{E}_{n}\,,b=0,1,2,3\}$. Denote by $\operatorname{wt}(E)$ the weight of an error $E$, which is the number of qubit positions on which an error $E\in{\mathcal{E}_{n}}$ acts non-trivially.
Given a projector onto a QECC subspace $\Pi$, the Knill-Laflamme conditions [34] states that the code has distance $\delta$, if and only if
@@
where $c_{ab}\in\mathds{C}$. A code is pure if $c_{ab}=\operatorname{tr}(E_{a}^{\dagger}E_{b})$ holds [9, 66]. Note that all codes with $K=1$ are defined to be pure. A code with $K=1$ is termed self-dual.
@@ 3.2. Stabilizer codes
Given $n-k$ commuting generators $g_{i}\in P_{n}$, let $S=\langle g_{1},\dots,g_{n-k}\rangle$ be a subgroup of the $n$-qubit Pauli group $P_{n}$ not containing $-\mathds{1}$. The stabilizer $S$ defines a code subspace, whose corresponding projector is given by
(9) $\Pi=\frac{1}{2^{n-k}}\prod^{n-k}_{i=0}(1+g_{i})=\frac{1}{2^{n-k}}\sum_{s_{i}\in S}s_{i}\,.$
In particular, the $s_{i}\in S$ stabilize $\Pi$, so that $s\Pi=\Pi s=\Pi$ for all $s\in S$.
A stabilizer code then encodes $k$ logical qubits into $n$ physical qubits with $n\geq k$ and therefore, $K=2^{k}$. The size of the stabilizer group is $|S|=\frac{2^{n}}{K}=2^{n-k}$ and when $k=0$ ($K=1$), we obtain a pure stabilizer state. For stabilizer codes, the coefficient $c_{ab}$ from Eq. (8) can only be ±$\pm 1$ or 00.
@@ 3.3. Quantum weight enumerators
(10) $A_{j}(\Pi)=$ $\sum_{\begin{subarray}{c}E\in{\mathcal{E}_{n}}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(E^{\dagger}\Pi)\operatorname{tr}(E\Pi)\,,$
(11) $B_{j}(\Pi)=$ $\sum_{\begin{subarray}{c}E\in{\mathcal{E}_{n}}\\ \operatorname{wt}(E)=j\end{subarray}}\operatorname{tr}(\Pi E^{\dagger}\Pi E)\,,$
where the sum runs over all errors $E$ of weight $j$ in the $n$-qubit Pauli basis ${\mathcal{E}_{n}}$. Let $K=\operatorname{rank}(\Pi)$, then
@@
for all $j=0,\dots,n$ and $A_{0}(\Pi)=K^{2}$. Self-dual codes ($K=1$) satisfy $A_{j}=B_{j}$ for all $j$. This is seen by expanding the definitions in Eq. (10) in terms of a projector onto a pure state $\Pi=|\psi\rangle\langle\psi|$.
@@
[55, Theorem 17] A projector $\Pi$ of rank $K$ on $n$ qubits corresponds to a quantum code of distance $\delta$, if and only if
@@
(14) $B(x,y)$ $=A\Big{(}\frac{x+3y}{2},\frac{x-y}{2}\Big{)}.$
where $A(x,y)$ is a polynomial defined as $A(x,y)=\sum^{n}_{j=0}A_{j}\,x^{n-j}y^{j}$ and likewise for $B(x,y)$. The coefficients of this polynomial transformation can be expanded as
@@
with $K_{j}(i):=K_{j}(i;n,4)$ being the quaternary Krawtchouk polynomial defined as [44],
(16) $K_{j}(i)=\sum^{i}_{\alpha=0}(-1)^{\alpha}3^{j-\alpha}\binom{i}{\alpha}\binom{n-i}{j-\alpha}\,.$
Note that, $KB_{0}(\Pi)=A_{0}(\Pi)$ corresponds to $\sum^{n}_{i=0}A_{i}(\Pi)=2^{n}K$. This implies that the weight enumerator of a self-dual code is invariant under the quantum MacWilliams transform.
The quantum shadow enumerator $S_{j}$ can be obtained through the transform [57]
(17) $S(x,y)$ $=A\Big{(}\frac{x+3y}{2},\frac{y-x}{2}\Big{)}\,.$
Its coefficients expand as
@@
It is known that $S_{j}\geq 0$ for all $j=0,\dots,n$, giving the so-called shadow bounds on quantum codes [56, 57]. If $K=1$ and $n-j$ odd then $S_{j}=0$ [57, Theorem 12].
In the case of qubit stabilizer codes, one has additionally the constraint that $\sum_{j\geq 0}A_{2j}=2^{n-\log_{2}(K)-1}$ or $\sum_{j\geq 0}A_{2j}=2^{n-\log_{2}(K)}$ for codes of type I and II respectively [9].
@@ 3.4. Linear programming bounds
The linear programming bounds are given by the conditions from Eqs. (3.3)–(18[67, 55],
find $A_{j},\,B_{j},\,S_{j}\geq 0\,,$
@@
$KB_{j}\geq A_{j}\quad\quad\quad\text{for }j=0,\dots,n\quad\text{with equality for $j<\delta$}\,,$
[…]
For given parameters $(\!(n,K,\delta)\!)_{2}$, if the linear program in Eq. (3.4) is infeasible, then a corresponding quantum codes does not exist.
In analogy to the classical case [13, 71], we will call the following subset of conditions the quantum Delsarte linear programming bounds for self-dual codes ($K=1$),
@@
$A_{j}\geq 0\ \,\,\text{with equality for }\,\,1<j<\delta\,,$
(20) $\sum^{n}_{i=0}K_{j}(i)A_{i}\geq 0\ \quad\text{for }j=0,\dots,n\,.$
If one has $\eta<2^{n}$ in Eq. (3.4), a code with parameters $(\!(n,1,\delta)\!)_{2}$ does not exist. We will formally derive this bound from a Lovász bound for self-dual quantum codes in Section 7.
@@
Note that for $K>1$ and without adding extra constraints like the quantum MacWilliams identity or the Knill-Laflamme conditions (that is $B_{j}=\sum_{i}K_{j}(i;n)A_{i}$ for $j=0,\dots,n$ and $KB_{j}=A_{j}$ for $j=0,\dots,\delta-1$) the quantum Delsarte bounds becomes weaker.
@@ 3.5. Lovász theta number of a graph
Let $G$ be a simple graph defined by a set of $N$ vertices $V$ connected by a set of edges $E$. Then the Lovász theta number can take the following two semidefinite programming formulations [20]. We call the following SDP1 for the Lovász number:
@@
$M_{ij}=0\quad\,\,\text{for all}\quad(i,j)\in E\,,$
@@
We call the following SDP2 for the Lovász number:
@@
$M_{ij}=0\quad\,\,\text{for all}\quad(i,j)\in E\,,$
@@
A slight variation of SDP1 from Eq. (3.5) with the additional entry-wise non-negativity constraint $M_{ij}\geq 0$ for all $i,j\in V$, defines $\vartheta^{\prime}(G)$. In the case of the Hamming cube, the computation of $\vartheta^{\prime}(G)$ can be shown to reduce to the classical Delsarte bound [64, 71],
@@
(23) $\sum_{i}K_{j}(i;n,2)x_{i}\geq 0\quad\text{for all }j=0,\dots,n\,.$
with $K_{j}(i;n,2)$ the binary Krawtchouk polynomial defined as [44],

### v2

@@ 2.2. Proof sketch: Lovász, Delsarte, and symmetry-reduced SDP bounds
    We consider a relaxation of the complete hierarchy of Result A that is indexed by $\langle E_{a}\rangle E_{a}$ with $E_{a}\in\mathcal{E}_{n}$. The key object is a matrix with entries $\Gamma_{ab}=\text{Re}({\langle E_{a}^{\dagger}\rangle}{\langle E_{b}\rangle}{\langle E_{a}E_{b}^{\dagger}\rangle})$. Such matrix is positive semidefinite by construction, and satisfies $\Gamma_{00}=1$, $\Gamma_{a0}=\Gamma_{aa}$, and $\Gamma_{ab}=0$ if $E_{a}E_{b}+E_{b}E_{a}=0$. For pure codes we additionally have $\Gamma_{ab}=0$ if $0<\operatorname{wt}(E_{a})<\delta$, $0<\operatorname{wt}(E_{b})<\delta$, or $0<\operatorname{wt}(E_{a}E^{\dagger}_{b})<\delta$.
@@
    Consider the case of self-dual codes, i.e., pure codes with $K=1$. Then the set of matrices $\Gamma$ with this structure can be identified with the Lovász theta body for a graph with loops, that is the feasible region of Eq. (3.5),
@@
    Consider the the anti-commutativity graph $G$ of the $n$-qubit Pauli group. Averaging $\mathrm{TH}(G)$ of under distance preserving automorphisms $\operatorname{Aut}(4,n)$ yields, along with additional equality constraints from projector conditions, the quantum Delsarte bound [Eq. (3.4)].
@@
    Using the Terwilliger algebra for $4$-ary codes, we average this main relaxation over distance-preserving automorphisms that keep the zero codeword invariant $\operatorname{Aut}_{0}(4,n)$. A block-diagonalization using the Terwilliger algebra [23] yields to a symmetry-reduced semidefinite program of size $O(n^{4})$.
A variation of the Lovász bound recovers the non-existence of the $(\!(7,1,4)\!)_{2}$ code, while a variation of the symmetry-reduced semidefinite programm shows the non-existence of $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes.
@@ 3. Key concepts
A quantum error-correcting code with parameters $(\!(n,K,\delta)\!)_{2}$ encodes a Hilbert space of dimension $K$ in $n$ qubits. A code of distance $\delta$ can correct all errors acting on at most $\lfloor\tfrac{\delta-1}{2}\rfloor$ qubits.
Define a $n$-qubit error basis (or Pauli basis) $\mathcal{E}_{n}$} by considering all $n$-fold tensor products of the identity and the three Pauli matrices,
(7) $I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\,,\quad X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,,\quad Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\,,\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\,$.
The elements of $\mathcal{E}_{n}$ then generate the $n$-qubit Pauli group $P_{n}=\{i^{b}E_{a}\,|\,E_{a}\in\mathcal{E}_{n}\,,b=0,1,2,3\}$. Denote by $\operatorname{wt}(E_{a})$ the Hamming weight of an error $E_{a}$, which is the number of positions an error $E_{a}\in{\mathcal{E}_{n}}$ acts on non-trivially.
Given a projector $\Pi$ onto a code subspace, the Knill-Laflamme conditions [35] state that the code has distance $\delta$, if and only if
@@
where $c_{ab}\in\mathds{C}$. A code is termed pure if $c_{ab}=\operatorname{tr}\!\left(E_{a}^{\dagger}E_{b}\Pi\right)/K=0$ for all $E_{a}^{\dagger}E_{b}\neq\mathds{1}$ with $0<\operatorname{wt}(E_{a}^{\dagger}E_{b})<\delta$ [9, 70]. Note that all codes with $K=1$ are assumed to be pure, as the Knill-Laflamme conditions would be trivial otherwise. Such codes are termed self-dual  [^4] This is an extension of the nomenclature for stabilizer codes, where $k=0$ ($K=1$) implies the stabilizer group is isomorphic to a self-dual additive classical code.. A motivation to study self-dual codes comes from the propagation rules for pure codes: any pure $(\!(n,K,\delta)\!)_{2}$ gives rise to a family of $(\!(n-s,2^{s}K,\delta-s)\!)_{2}$ codes for $s=1,\dots,\delta-1$ [59, Theorem 19]. An additional interest lies in the fact that self-dual codes correspond to absolutely maximally entangled and $m$-uniform states [64, 29].
@@ 3.2. Stabilizer codes
Given nkn-k commuting independent generators $g_{i}\in P_{n}$, let $S=\langle g_{1},\dots,g_{n-k}\rangle$ be a subgroup of the $n$-qubit Pauli group $P_{n}$ not containing $\mathds{1}$. The stabilizer group $S$ defines a code subspace, whose corresponding projector is,
(9) $\Pi=\frac{1}{2^{n-k}}\prod^{n-k}_{i=1}(\mathds{1}+g_{i})=\frac{1}{2^{n-k}}\sum_{s\in S}s\,.$
In particular, the $s_{i}\in S$ stabilize $\Pi$, so that $s\Pi=\Pi s=\Pi$ for all $s\in S$. A stabilizer code encodes $k$ logical qubits into $n$ physical qubits with $k\leq n$ and $K=2^{k}$. The size of the stabilizer group is $|S|=\frac{2^{n}}{K}=2^{n-k}$ and when $k=0$ ($K=1$) one has a stabilizer state.
@@ 3.3. Quantum weight enumerators
(10) $A_{j}(\Pi)=$ $\sum_{\begin{subarray}{c}E_{a}\in{\mathcal{E}_{n}}\\ \operatorname{wt}(E_{a})=j\end{subarray}}\operatorname{tr}(E_{a}^{\dagger}\Pi)\operatorname{tr}(E_{a}\Pi)\,$,
(11) $B_{j}(\Pi)=$ $\sum_{\begin{subarray}{c}E_{a}\in{\mathcal{E}_{n}}\\ \operatorname{wt}(E_{a})=j\end{subarray}}\operatorname{tr}(E_{a}\Pi E_{a}^{\dagger}\Pi)\,$,
where the sum runs over all errors $E_{a}$ of weight $j$ in the $n$-qubit Pauli basis $\mathcal{E}_{n}$}. Let $K=\operatorname{rank}(\Pi)$, then we have
$A_{0}(\Pi)$ $=K^{2}\,$,
@@
for all $j=0,\dots,n$. Self-dual codes satisfy $A_{j}(\Pi)=B_{j}(\Pi)$ for all $j$. This is seen by expanding $A_{j}$ and $B_{j}$ in terms of a projector onto a pure state $\Pi=|\psi\rangle\langle\psi|$.
@@
[59, Theorem 17] A projector $\Pi$ of rank $K$ on $n$ qubits corresponds to a quantum code of distance $\delta$, if and only if
@@
(14) $B(x,y)$ $=A\Big{(}\frac{x+3y}{2},\frac{x-y}{2}\Big{)}\,$,
where $A(x,y)$ is a polynomial defined as $A(x,y)=\sum^{n}_{j=0}A_{j}\,x^{n-j}y^{j}$ and likewise for $B(x,y)$ 666 Technically speaking, the $A(x,y)$ is the weight enumerator, whose coefficients $A_{j}$ form the weight distribution. We use the notions interchangeably.. The coefficients of this polynomial transformation can be expanded as
@@
where $K_{j}(i;n):=K_{j}(i;n,4)$ are the quaternary Krawtchouk polynomials [45], defined as
(16) $K_{j}(i;n)=\sum^{n}_{\alpha=0}(-1)^{\alpha}3^{j-\alpha}\binom{i}{\alpha}\binom{n-i}{j-\alpha}\,$.
Note that $KB_{0}(\Pi)=A_{0}(\Pi)=K^{2}$ corresponds to $\sum^{n}_{i=0}A_{i}(\Pi)=2^{n}K$. The weight enumerator of a self-dual code is invariant under the quantum MacWilliams transform. In the case of qubit stabilizer codes, one has the additional constraint that $\sum_{j=0}^{\lfloor n/2\rfloor}A_{2j}(\Pi)=2^{n-\log_{2}(K)-1}$ or $\sum_{j=0}^{\lfloor n/2\rfloor}A_{2j}(\Pi)=2^{n-\log_{2}(K)}$ for codes of type I and II respectively [9, 70], as well as the constraint that the weight enumerators must take integer values.
The quantum shadow enumerator $S_{j}$ can be obtained through the polynomial transform [62]
(17) $S(x,y)$ $=A\Big{(}\frac{x+3y}{2},\frac{y-x}{2}\Big{)}\,$,
and its coefficients expand as
@@
It is known that $S_{j}(\Pi)\geq 0$ for all $j=0,\dots,n$, giving the so-called shadow bounds on quantum codes [60, 62]. If $K=1$ and njn-j is odd then $S_{j}(\Pi)=0$ [62, Theorem 12].
@@ 3.4. Linear programming bounds
The enhanced linear programming bounds are given by the conditions from Eqs. (3.3)–(18[71, 59],
find $A_{j},\,B_{j},\,S_{j}\geq 0$
@@
(19) $KB_{j}\geq A_{j}\quad\quad\text{with equality for }0\leq j<\delta\,$,
where $A_{j}$, $B_{j}$, $S_{j}$ are related by Eqs. (14) and (17). In case of self-dual, pure, or stabilizer codes the above-mentioned extra constraints apply. For given parameters $(\!(n,K,\delta)\!)_{2}$, if the linear program in Eq. (3.4) is infeasible, then a corresponding quantum codes does not exist.
In analogy to the classical case (see Eq. (3.5) and [13, 74]), for self-dual codes we will call the following the quantum Delsarte linear programming bound,
@@
$A_{j}\geq 0\quad\quad\text{with equality for }0<j<\delta\,$,
(20) $\sum^{n}_{i=0}K_{j}(i;n)A_{i}\geq 0\quad\text{for all }j=0,\dots,n\,.$
From $\sum_{i=0}^{n}A_{i}(\Pi)=2^{n}$, it follows that if $\eta<2^{n}$, then a code with parameters $(\!(n,1,\delta)\!)_{2}$ does not exist. It is clear that Eq. (3.4) relaxes the constraints in Eq. (3.4). As we show in Section 7, it can also be obtained by symmetrization of a Lovász bound for self-dual quantum codes. This mirrors the derivation of the classical Delsarte bound [Eq. (3.5)].
@@
Note that for $K>1$ and without adding extra constraints such as the quantum MacWilliams identity or the Knill-Laflamme conditions (i.e., $B_{j}=\sum_{i=0}^{n}K_{j}(i;n)A_{i}$ for $j=0,\dots,n$ and $KB_{j}=A_{j}$ for $j=0,\dots,\delta-1$) the quantum Delsarte bounds becomes weaker.
@@ 3.5. Lovász theta number of a graph
Let $G$ be a simple graph defined by a set of $N$ vertices $V$ connected by a set of edges $E$. For the Lovász theta number there exist a number of equivalent definitions. We consider the following two semidefinite programming formulations [20]. We call the following SDP1 for the Lovász theta number:
@@
$M_{ij}=0\quad\,\,\text{for all}\quad\{i,j\}\in E\,$,
@@
We call the following SDP2 for the Lovász theta number:
@@
$M_{ij}=0\quad\,\,\text{for all}\quad\{i,j\}\in E\,$,
@@
A slight variation of SDP1 from Eq. (3.5) with the additional entry-wise non-negativity constraint $M_{ij}\geq 0$ for all $i,j\in V$, defines $\vartheta^{\prime}(G)$. For binary codes, the computation of $\vartheta^{\prime}(G)$ can be shown to reduce to the classical Delsarte bound [68, 74],
@@
(23) $\sum_{i=0}^{n}K_{j}(i;n,2)x_{i}\geq 0\quad\text{for all }j=0,\dots,n\,.$
where $K_{j}(i;n,2)$ are the binary Krawtchouk polynomials defined as [45],